### v1

# ReCap: Better Gaussian Relighting with Cross-Environment Captures

Abstract
Accurate 3D objects relighting in diverse unseen environments is crucial for realistic virtual object placement. Due to the albedo-lighting ambiguity, existing methods often fall short in producing faithful relights. Without proper constraints, observed training views can be explained by numerous combinations of lighting and material attributes, lacking physical correspondence with the actual environment maps used for relighting. In this work, we present ReCap, treating cross-environment captures as multi-task target to provide the missing supervision that cuts through the entanglement. Specifically, ReCap jointly optimizes multiple lighting representations that share a common set of material attributes. This naturally harmonizes a coherent set of lighting representations around the mutual material attributes, exploiting commonalities and differences across varied object appearances. Such coherence enables physically sound lighting reconstruction and robust material estimation — both essential for accurate relighting. Together with a streamlined shading function and effective post-processing, ReCap outperforms the leading competitor by $3.4$ dB in PSNR on an expanded relighting benchmark.
Figure 1: ReCap shows a significant advantage in reconstructing environment maps with accurate tones and color fidelity. Both qualitative and quantitative assessments show that ReCap achieves more realistic and consistent relighting results under a range of unseen lighting conditions. See the project page.
1 Introduction
For a realistic and immersive augmented reality experience, virtually placed objects must convincingly reflect light, cast shadows, and adapt naturally to different lighting conditions. Achieving the desired realism requires a physically accurate response to environment lighting, driving a line of research at enabling relighting capabilities in popular neural representation models.
In recent years, Neural Radiance Field (NeRF) [27] gained prevailing popularity as an implicit scene representation. Subsequent NeRF-based relighting methods [37, 48, 19] produced impressive relighting results, but their computational demands make them impractical for interactive applications [18, 44]. Lately, 3D Gaussian Splatting (3DGS) [22] is widely acclaimed as an explicit 3D representation model for its high rendering quality and interactive frame rates, naturally suited for applications requiring real-time performance. Building on this, follow-up works [18, 25, 44, 15] have enabled relighting of Gaussians using explicit shading functions and learnable lighting representations, often in the form of environment maps.
While standard HDR maps could theoretically replace these learned environment maps for relighting, directly substituting them, as current methods do, remains questionable due to the unclear physical meaning of the learned values. Because of the albedo-lighting ambiguity [1, 5], where changes in surface albedo are indistinguishable from changes in lighting intensity, existing supervision from the reconstruction loss alone is not enough for a truthful lighting reconstruction. As shown in Fig. 1a), the learned environments are often observed to be tinted with object colors, shifted in tone, scaled in intensity or filled with noise. Without proper constraints, these maps act as sinks for unmodeled residual terms during optimization, becoming indispensable for producing high-quality novel view synthesis results. Replacing them with the ground truth HDR maps significantly degrades the output quality, much less when attempting to relight with novel environment maps.
Inspired by photometric appearance modeling [32, 24, 43] which estimates surface properties from object appearances under varied lighting, we introduce additional photometric supervision to address the albedo-lighting ambiguity. While traditional approaches rely on controlled lighting setups like light stages [11] and collocated lights [42] to provide supervision through known light directions and/or intensities, such controlled capture is impractical for real-world applications. Instead, we propose ReCap to leverage object captures across unknown lighting conditions, modeling light-dependent appearances with multiple environment maps that share a common Gaussian model. Conceptually, this resembles multi-task learning, where the learned environment maps act as task heads querying a shared material representation for varied object appearances. In this case, the “querying” is done by the physically-based shading function, which facilitates accurate separation of material and lighting. This joint optimization fosters robust material estimation across diverse environments while maintaining internal consistency and coherence, simultaneously supporting precise lighting reconstruction.
To streamline the joint optimization, we introduce a generalized shading function based on the split-sum approximation [21], which eliminates a material parameter that introduces ambiguity in inverse rendering. Additionally, we ensure compatibility with standard HDR maps by encouraging the learned values to remain in a linear HDR space through appropriate post-processing. This enables the direct application of novel environment maps without the need for image normalization [26] or map adjustments [25].
Current relighting evaluations are limited in scope, focusing primarily on diffuse surfaces [19, 49]. For a more comprehensive assessment, we re-rendered $13$ objects from NeRF [27] and RefNeRF [38] featuring both diffuse and specular surfaces. As shown in Fig. 1, experiments on the expanded benchmark confirm the effectiveness of ReCap in producing robust and more realistic relighting results. Codes and datasets will be made publicly available here.
To summarize, the contribution of this work includes
  • Treating cross-environment captures as multi-task targets, we address the albedo-lighting ambiguity via the joint optimization of shared material properties and independent lighting representations.
  • We propose a novel shading function with physically appropriate post-processing, providing more flexible material representation that eases optimization and allows the direct application of standard HDR maps.
  • ReCap achieves state-of-the-art relighting performance on a more comprehensive benchmark, showing robustness across diverse lighting and object types.
Figure 2: The proposed ReCap training framework. Compared to original 3DGS [22], each Gaussian is augmented with 3 extra material attributes. Given $k$ sets of object appearances from unknown lighting conditions as input, $k$ learnable environment maps are instantiated. Gaussian color is pre-computed according to the shading function in the world space based on environment queries and material properties. 2D images are rasterized with standard Gaussian splatting and used for loss computation. $\mathcal{L}_{\text{image}}$: image reconstruction loss from [22]. Additional loss terms for material and geometry are not shown.
2 Related Work
Novel View Synthesis (NVS).
Pure NVS techniques focus on reproducing the scene appearance under its original environment. NeRF [27] and its follow-ups [2, 3, 4] achieve remarkable NVS quality with implicit radiance fields and volume rendering. Although great progress has been made in NeRF acceleration [14, 28, 41] leveraging voxel grids or hash tables, their rendering speeds ($\sim$10 fps) are still far from interactive. Building on progress in differentiable point-based rendering [45, 23], 3DGS [22] recently introduced an explicit Gaussian representation with an efficient tile-based rasterizer, delivering photo-realistic rendering at impressive frame rates ($\sim$100 fps). This makes 3DGS well-suited for high-quality real-time relighting.
Relighting.
The relighting task traditionally involves editing lighting within fixed views [12, 35, 42]. The term has since broadened to encompass novel view relighting [43, 20], which aims to generate images from both new viewpoints and under new lighting conditions. With the advent of differentiable neural rendering, reconstruction-based methods [7, 46, 48, 37] have found a promising direction to combine physically-based rendering (PBR) with neural representations for editable and realistic rendering. Closely related to our work, PBR is also incorporated by 3D Gaussians, which enable Gaussian relighting. 3DGS-DR [44] and GShader [18] both specialize in specular objects and propose customized shading functions. R3DG [15] further incorporates ray tracing for indirect illumination, while GS-IR [25] relies on baked occlusion maps. With only single environment captures as input, these Gaussian relighting methods face challenges in correctly decoupling lighting from material properties.
Albedo-lighting Disambiguation.
Recovering reliable illumination and albedo from object appearances is an ill-posed problem [34] in inverse rendering. It is often simplified under the assumption of controllable lighting [6, 30], known geometry [31], or when regularized with strong priors [36, 13] in domain-specific tasks such as relighting of human faces. For more general cases, object captures from varied lighting, known as photometric images, can provide important visual cues. Several NeRF-based relighting methods have taken advantage of such captures. NeRV [37] requires multiple known lighting environments, while TensoIR [19] internalizes lighting as an additional dimension in their factorized tensorial representation. NeRD [7] models lighting as spherical Gaussians and optimizes separate lighting for every input image. In contrast, we use object captures from unknown lighting conditions, with learnable lighting representations optimized independently for each scene as environment maps.
3 Method
In this section, we introduce ReCap, a robust Gaussian relighting method that leverages cross-environment object captures. The overall framework is illustrated in Fig. 2. In Sec. 3.1, we explain how relighting is enabled in 3DGS with an explicit shading function. In Sec. 3.2, an optimization-friendly variant of the split-sum-approximation is introduced for shading. Lighting representation and post-processing are detailed in Sec. 3.3 and Sec. 3.4, with normal estimation covered in Sec. 3.5.
3.1 Relighting Gaussians with shading function
3DGS represents objects as explicit point clouds. A point at location $\mathbf{x}$ in the world space is represented as a 3D Gaussian function defined by covariance matrix $\Sigma$ and mean $\mu$:
$\mathcal{G}(\mathbf{x}|\mu,\Sigma)=e^{\frac{1}{2}(\mathbf{x}-\mu)^{T}\Sigma^{-1}(\mathbf{x}-\mu)}$ (1)
Additionally, each point holds an opacity attribute $\alpha$ and a color attribute $\mathbf{c}$ for point-based $\alpha$-blending. From a viewing direction $\mathbf{v}$, the pixel color $C$ can be computed by blending ordered points overlapping the pixel:
$C(\mathbf{v})=\sum_{i}\mathbf{c}_{i}(\mathbf{v})\alpha_{i}\prod_{j=1}^{j-1}(1-\alpha_{j})$ (2)
The original 3DGS models the view-dependent color $\mathbf{c}(\mathbf{v})$ with spherical harmonics. This simplification abstracts the complex view-dependent interactions between material, lighting and geometry into a composite representation. Relighting Gaussians thus requires an alternative representation of $\mathbf{c}(\mathbf{v})$ that factors out the influence of lighting. A natural choice is to use various well-established shading functions from graphics. Ignoring any emission, let the classic rendering equation represent the outgoing radiance at Gaussian point $\mathbf{x}$ viewed from direction $\mathbf{v}$ as
[…]
where $\Omega$ is the hemisphere above the surface, $f_{r}$ is the bi-directional reflection function (BRDF), $\mathbf{l}$ is the incident light direction and $L_{in}$ is the incoming radiance, all defined in the local coordinate system centered at $\mathbf{x}$. The Gaussian color $\mathbf{c}$ can be computed from $L_{out}$ with proper post-processing such as tone mapping and gamma correction as discussed in Sec. 3.4. The pixel color $C$ becomes
$C(\mathbf{v})=\sum_{i}\mathbf{c}_{i}(\mathbf{v}|f_{r},L_{in},\mathbf{n})\alpha_{i}\prod_{j=1}^{j-1}(1-\alpha_{j}).$ (4)
The rendering equation is commonly simplified for implementation as shading functions. GShader [18] and GS-DR [44] handcrafted their shading functions with a light-independent diffuse component and a light-dependent specular component querying lighting information from a learned environment map. R3DG [15] and GS-IR [25] both leverage microfacet based model as a more expressive alternative. In our work, we start with the split-sum approximation of the microfacet model and propose a more optimization-friendly variant.
3.2 Disambiguate Split Sum Approximation
Disney Principled BRDF [8] provides user-friendly parameters building on the Cook-Torrance microfacet BRDF [9]. We adopt simplifications from Epic Game [21] and consider the following parameters: 1) basecolor, $\mathbf{b}\in[0,1]^{3}$; 2) roughness, $r\in[0,1]$; 3) metallic, $m\in[0,1]$ ; and 4) specular, $s=0.04$, assumed to be constant for non-metals.
The BRDF of interest is given as
$\begin{split}f_{r}(\mathbf{x},\mathbf{v},\mathbf{l})=&(1-m)\frac{\mathbf{b}}{\pi}\\ &+\frac{D(r,\mathbf{n},\mathbf{l},\mathbf{v})F(\mathbf{b},m,s,\mathbf{l},\mathbf{v})G(r,\mathbf{n},\mathbf{l},\mathbf{v})}{4(\mathbf{l}\cdot\mathbf{n})(\mathbf{v}\cdot\mathbf{n})},\end{split}$ (5)
where $D$, $F$ and $G$ are the normal distribution function, the Fresnel term and the geometry term respectively.
Substitute Eq. 5 into Eq. 3 and apply the split-sum approximation as described in  [21], the shading function can be written as
$\begin{split}L_{out}(\mathbf{x},\mathbf{v})&=\underbrace{E_{d}(\mathbf{n})(1-m)\mathbf{b}}_{diffuse}\\ &+\underbrace{E_{s}(\mathbf{n},\mathbf{v})\left[F_{0}\beta_{1}(r,\mathbf{n},\mathbf{v})+\beta_{2}(r,\mathbf{n},\mathbf{v})\right]}_{specular}\\ \end{split}$ (6)
where $E_{d}$ and $E_{s}$ are the pre-filtered environment maps for diffuse and specular reflectance, $\beta_{1}$ and $\beta_{2}$ are pre-calculated BRDF look-ups, $F_{0}=m\mathbf{b}+(1-m)s$ is the effective reflectance.
This is a linear blend of two different models controlled by the metallic parameter. For metals, there is no diffuse component and the specular part is colored. For non-metals, the diffuse part is colored but not the specular part. Dropping arguments for conciseness, the two models are
$\begin{split}L_{\text{metal}}&=E_{s}\mathbf{b}\beta_{1}+E_{s}\beta_{2}\\ L_{\text{non-metal}}&=E_{s}s\beta_{1}+E_{s}\beta_{2}+E_{d}\mathbf{b}.\\ \end{split}$ (7)
Without prior knowledge on the distribution of material parameters of real metal and non-metal, optimization of the metallic parameter is problematic especially when the lighting is also learned. First, $s$ and $\mathbf{b}$ have overlaps. Certain specular highlights of dark surfaces may be misinterpreted and assigned to the wrong metallic value, as illustrated by the example in Fig. 3. Furthermore, when the environment maps and roughness values are optimized such that $E_{d}\sim E_{s}\beta_{1}$, the two equations becomes interchangeable with some value of $\mathbf{b}$ leading to ambiguous optimization.
Instead, we discard the metallic parameter and expand the specular parameter to propose a general expression as
$L_{\text{out}}=E_{s}\mathbf{s}\beta_{1}+E_{s}\beta_{2}+E_{d}\mathbf{b},$ (8)
where $\mathbf{s}\in[0,1]^{3}$ is now a vector representing specular tint. For $\mathbf{s}=[s,s,s]^{T}$, it becomes $L_{\text{non-metal}}$; for $\mathbf{s}=\mathbf{b}_{\text{metal}}$ and $\mathbf{b}=0$, it becomes $L_{\text{metal}}$. On the downside, the expanded range of $\mathbf{s}$ encompasses a significant portion of unnatural specular colors. To avoid overly saturated specular tint, we apply saturation penalty on $\mathbf{s}$ as
$\mathcal{L}_{\text{sat}}=\lambda_{\text{sat}}\cdot\left\|\mathbf{s}-\mathbf{s}_{\text{mean}}\right\|.$ (9)
To account for energy conservation, we further include a regularizer to encourage $\left\|\mathbf{s}\right\|+\left\|\mathbf{b}\right\|\leq 1$.
Figure 3: With the original shading model, the shield of the helmet is falsely identified as being metallic during optimization.
3.3 Lighting Representation
While spherical functions [39, 46, 16] are popular choices for efficient lighting representations, we adopt an image based lighting [10] model to provide high frequency details necessary for specular reflections. Specifically, each environment lighting is represented by a $6\times 256\times 256$ learnable cube map, which is pre-filtered into a diffuse map, $E_{d}$, for diffuse reflection [33] and a set of specular mipmaps across different roughness levels, $E_{s}$, for specular reflection [21]. In practice, we use the efficient approximation provided by NVDIFFRAST [29] for differentiable pre-filtering and querying of the learnable environment maps, which are performed in every forward pass for shading.
To leverage photometric supervision from cross-environment captures, $k$ sets of learnable environment maps will be instantiated to explain $k$ sets for object appearances. As illustrated by Fig. 4, the same object position can display a range of pixel colors depending on viewing direction and lighting. With only single-environment captures, the view-dependent variations can be explained by a multitude of material-lighting combinations, manifesting the albedo-lighting ambiguity. By introducing additional sets of object appearances under new, unknown environments, the light-dependent variations are attributed to corresponding learnable environment maps as multi-task learning targets. The joint optimization encourages physically sound decoupling of material properties and lighting to explain all observed appearances as guided by the shading function.
While increasing the number of environments can enhance the decoupling as later shown in Sec. 4.3, it also requires additional effort to capture real objects across multiple scenes. As a proof of concept, we limit our investigation to a dual-environment setup for the remainder of the text.
Figure 4: The same object position exhibit view-dependent and light-dependent pixel color, the later is accounted for by querying corresponding learnable environments.
3.4 Physically Appropriate Post-processing
Standard HDR maps typically represent radiance in linear RGB space when used as lighting sources in rendering [10]. To properly apply them for relighting, the learned lighting values need a clear physical interpretation. Current methods make diverse design choices during training, which jointly determine the interpretation of the learned environment maps and how new HDR maps should be processed. For example, if the range of the learned values is limited between $[0,1]$ and gamma correction is applied after the shading equation, the learned environment map is interpreted as being in a linear low dynamic range (LDR) space. In Tab. 1, we analyze three major factors that impact relighting performance. Gamma correction is found to be essential for relighting, which is expected, since the shading function models linear light transport. However, it is often overlooked [15, 44], as its omission still provides good NVS results. Complex tone-mappers are common in graphics and are widely adopted [37], but they perform worse than simple clipping, likely due to the introduced non-linearity that hinders optimization.
Striking a balance between relighting and NVS, we constrain our environment maps to be positive, applying simple clipping followed by gamma correction. This ensures a linear HDR lighting space, where new environment maps can be applied without modification. Together with cross-environment decoupling, our learned environment maps are closer to the actual HDR maps in distribution. This removes the need for normalizing relit images [26] to match the average colors of ground-truth images (which are not available in practice) when comparing relighting results.
Table 1: Relighting and NVS performance (PSNR) of various design choices. The $\rightarrow$ symbol denotes implication on the learned environment map. During relighting, the input HDR map are pre-processed to match the interpretation. See the supplementary for more details.
Env. Map Range Tonemap Gamma Relight NVS
[…]
3.5 Geometry Estimation
Accurate normal estimation is essential for precise querying of the high frequency light maps. As discrete sparse point clouds, 3D Gaussians have no native support for normal computation. Following the common observation that converged Gaussians are often flat and align with the object surface [25, 18, 17, 44], we adopt the shortest axis of each Gaussian as an estimation of the normal vector. This shortest-axis normal approximation, $\mathbf{n}$, is simply constrained using depth-derived normal, $\mathbf{\hat{n}}$, where the depth image is rendered from Gaussian opacities. This depth-normal consistency loss is given as
$\mathcal{L}_{\text{dn}}=\lambda_{\text{dn}}\cdot\left\|\mathbf{n}-\mathbf{\hat{n}}\right\|^{2}.$ (10)
While existing works commonly augment normal estimation with additional residuals [18] or learn per-Gaussian normal vectors directly [15, 25], we observe that such learnable normals often overfit to explain specularities or shadowing effects. This leads to improved NVS but compromises geometric accuracy. As shown in Fig. 5, the cross-lighting consistency imposed by ReCap naturally improves normal estimation by preventing overfitting to a single lighting condition. When only single-environment captures are used, specular highlights become embedded in the surface normals, preserving the specific highlight shapes seen in training even when relighted with new environment maps. In contrast, our ReCap with dual lighting produce accurate highlight shapes with better surface normal.
Figure 5: The comparison of estimated normal and corresponding relighting results. With single-environment captures, the highlights from the train view are falsely attributed to object property instead of lighting, passing down to relighting views. Cross-environment supervision provides more robust normal estimation and correct highlight shapes.
Table 2: Relighting results across different environment maps. ReCap trains on 100 views each from “courtyard” and “bridge”, while other methods use 200 views from “bridge” only. The * symbol indicates learned environment, equivalent to NVS. Note that 3DGS-DR’s good performance on “bridge” reflects its tendency to regurgitate training views. Best results are highlighted in bold.
Unseen Envs Training Envs NVS
night snow sunset interior fireplace forest average courtyard bridge bridge*
[…]
GShader [18] 22.33 18.14 19.26 21.87 21.03 23.04 20.95 20.68 26.04 32.92
3DGS-DR [44] 20.88 18.92 23.53 21.00 19.52 25.19 21.51 20.58 33.19 35.50
GS-IR [25] 24.87 21.95 21.15 20.18 24.27 20.47 22.15 20.02 20.72 29.95
R3DG [15] 23.20 22.05 22.93 19.65 22.62 22.26 22.12 20.22 22.25 30.44
ReCap (ours) 26.68 24.56 26.56 24.34 25.51 25.45 25.52 25.51 26.97 32.03
[…]
GShader [18] 0.899 0.889 0.900 0.907 0.869 0.907 0.895 0.893 0.950 0.969
3DGS-DR [44] 0.892 0.897 0.931 0.909 0.858 0.910 0.900 0.894 0.971 0.978
GS-IR [25] 0.900 0.896 0.892 0.878 0.876 0.863 0.884 0.871 0.879 0.938
R3DG [15] 0.889 0.914 0.920 0.865 0.867 0.879 0.889 0.871 0.879 0.959
ReCap (ours) 0.932 0.938 0.951 0.929 0.903 0.927 0.930 0.942 0.951 0.965
[…]
GShader [18] 0.086 0.115 0.095 0.091 0.106 0.089 0.102 0.097 0.061 0.043
3DGS-DR [44] 0.077 0.089 0.070 0.086 0.101 0.081 0.084 0.095 0.040 0.033
GS-IR [25] 0.104 0.102 0.105 0.112 0.112 0.114 0.108 0.122 0.113 0.071
R3DG [15] 0.104 0.088 0.081 0.126 0.113 0.103 0.108 0.113 0.113 0.053
ReCap (ours) 0.062 0.070 0.057 0.076 0.076 0.071 0.069 0.065 0.058 0.047
4 Experiments
4.1 Implementation Details
Datasets.
We construct a more comprehensive RelightObj dataset by relighting 8 general objects from NeRF Synthetic Dataset [27] and 5 shiny objects from the Shiny Blender Dataset [38] under 8 different HDR maps. Each scene includes 200 training views and 200 test views, with identical camera poses across scenes. For a consistent relighting comparison, all methods are trained with captures from a single environment, except ReCap, which additionally uses a second environment. To ensure fairness in the number of training views, all methods use 200 views from the primary training environment, while ReCap uses 100 views from each of the two training environments. Relighting performance will be judged on unseen environments.
Baseline and Metrics.
We compare our method against the following baselines: a) GShader [18]: A Gasussian rendering method utilizing a customized shading function with a residual color term. b) GS-IR [25]: An inverse rendering approach using view-space shading with a microfacet BRDF model and a baked occlusion map. c) 3DGS-DR [44]: A Gasussian rendering method leveraging a customized shading function and normal propagation in view-space. d) R3DG [15]: A Gaussian relighting technique that incorporates ray tracing and a microfacet BRDF model. Following these works, rendering and relighting quality are evaluated using PSNR, SSIM [40], and LPIPS [47].
Training and Testing.
All experiments are conducted on an Nvidia RTX 3090 graphics card. We optimize our models using the Adam optimizer over 30,000 iterations. All methods are retrained on our dataset using their official repositories with recommended settings, results are reported from the retrained models. Since GShader [18] and 3DGS-DR [44] do not provide official relighting code, we pre-processed the HDR maps according to their design choices, with details in the supplementary material.
Figure 6: Qualitative comparison of relighting results across different environment maps. Specular objects provide more information for lighting reconstruction, leading to better albedo-lighting decoupling hence more truthful relighting results. Refer to the supplementary for more visuals.
Table 3: Object-wise relighting performance. The average is reported over unseen environment maps. ReCap shows the best overall performance for both diffuse and specular objects, while performing comparably on objects with shadows or translucent materials. Best results are highlighted in bold.
Relit NeRF Synthetic Relit Shiny Blender
chair drums ficus hotdog lego ship material mic helmet coffee musclecar toaster teapot
[…]
GShader [18] $20.83$ $19.69$ $25.56$ $18.80$ $18.85$ $17.10$ $20.03$ $27.86$ $19.07$ $\mathbf{20.28}$ $21.51$ $18.55$ $24.14$
3DGS-DR [44] $21.25$ $21.82$ $25.78$ $19.34$ $24.37$ $18.98$ $20.01$ $28.02$ $20.33$ $18.56$ $19.47$ $17.35$ $24.34$
GS-IR [25] $21.87$ $22.37$ $26.75$ $21.74$ $19.89$ $\mathbf{21.18}$ $22.26$ $27.25$ $19.31$ $18.60$ $23.39$ 16.79 26.52
R3DG [15] 23.67 21.23 25.60 24.38 22.21 20.50 21.55 27.67 21.17 16.27 22.74 16.66 23.89
ReCap (ours) 26.63 24.77 28.24 23.25 24.44 20.12 25.83 31.44 23.55 19.78 29.14 25.05 29.49
[…]
GShader [18] 0.914 0.891 0.951 0.885 0.866 0.774 0.871 0.968 0.889 0.931 0.898 0.824 0.973
3DGS-DR [44] 0.920 0.911 0.958 0.882 0.926 0.771 0.875 0.970 0.885 0.921 0.881 0.824 0.971
GS-IR [25] 0.905 0.901 0.951 0.886 0.856 0.783 0.873 0.958 0.850 0.896 0.896 0.774 0.967
R3DG [15] 0.924 0.895 0.952 0.904 0.888 0.775 0.864 0.962 0.865 0.877 0.900 0.784 0.968
ReCap (ours) 0.951 0.937 0.972 0.912 0.918 0.817 0.921 0.978 0.920 0.928 0.944 0.910 0.983
[…]
GShader [18] 0.078 0.068 0.038 0.133 0.099 0.187 0.089 0.024 0.134 0.142 0.073 0.157 0.038
3DGS-DR [44] 0.055 0.056 0.029 0.110 0.055 0.173 0.081 0.020 0.116 0.144 0.075 0.143 0.033
GS-IR [25] 0.068 0.069 0.040 0.133 0.096 0.189 0.093 0.032 0.178 0.164 0.083 0.220 0.041
R3DG [15] 0.060 0.068 0.037 0.105 0.084 0.188 0.103 0.033 0.171 0.157 0.078 0.210 0.038
ReCap (ours) 0.042 0.042 0.023 0.091 0.065 0.167 0.063 0.016 0.094 0.128 0.047 0.088 0.027
4.2 Comparison with Previous Methods
Scene-wise Comparison.
ReCap consistently outperforms other methods on all three metrics across various unseen environments, as shown in  Tab. 2, while maintaining comparable NVS performance. As further supported by the visual comparisons in Fig. 6 where ReCap demonstrates noticeable perceptual improvements over other methods, this indicates advantageous relighting quality and fidelity without sacrificing NVS performance.
A notable observation across all methods is the large performance gap between NVS using the learned environment map and relighting using the ground truth HDR map (bridge* vs bridge). This highlights the albedo-lighting ambiguity, where the learned lighting representation diverges from ground truth but still works effectively with the learned materials for NVS. Among all methods, 3DGS-DR [44] shows the smallest gap, nonetheless its performance on unseen environments remains limited. As shown by the second row in Fig. 6, this is because that 3DGS-DR tends to preserve a consistent tone across different environments, closely matching the training views and adapting only the specular reflections based on new lighting.
Object-wise Comparison.
While scene-wise comparison confirms our method’s ability to generalize across different lighting conditions, a more granular evaluation of how effectively it adapts to diverse object attributes is presented in Tab. 3. Across various test objects featuring diffuse and specular surfaces, ReCap shows the best overall performance. Especially with more specular objects (e.g., toaster, musclecar) where other methods find challenging, ReCap shows a unique advantage in decoupling rich reflected lighting details for accurate relighting.
Reconstructed Environments.
While specular objects pose significant challenges for NVS due to their complex view-dependent appearance[18, 44], they offer distinct advantages for relighting tasks. Specular surfaces act as natural light probes, providing abundant high-frequency, view-dependent details that facilitate more accurate environment map reconstruction. In Fig. 7, we display the reconstructed environment maps for objects with varying level of specularity. When the object is mostly diffuse, it’s challenging to disentangle lighting effects from the object’s intrinsic properties due to insufficient lighting cues.
Figure 7: Learned environment maps for different objects. As objects become more diffuse, reconstructions are less accurate.
4.3 Ablation Study
Shading function.
To demonstrate the effectiveness of our proposed shading function in Eq. 8, we compare the average relighting performance across unseen scenes in  Tab. 4. The new shading function yields the most substantial improvement by offering a more flexible material representation, with the two additional regularization terms also contributing positively to the final results. Additional visual examples are provided in the supplementary material.
Table 4: Ablation of shading function and regularization terms. $\mathcal{L}_{sat}$: specular saturation loss. $\mathcal{L}_{ec}$: energy conservation loss.
Proposed Shading $\mathcal{L}_{sat}$ $\mathcal{L}_{ec}$ PSNR SSIM LPIPS
- - - 24.62 0.923 0.076
- - 25.29 0.929 0.070
- 25.38 0.929 0.069
25.52 0.930 0.069
Extra training environments.
Up to this point, we have focused on a dual-environment setup. In this experiment, we examine how extra photometric supervision from additional training environments affects the relighting results. Due to the constraint of 200 training views per scene, we allocate 100 identical views per environment as we include more environments for training. To provide a basis for comparison with existing results, we additionally compare the use of identical views versus extra views.
Results in Tab. 5 highlight three key findings: 1) more views from additional environment yields better results than adding more views within the same environment; 2) using identical views across environments provides stronger feature decoupling; and 3) increasing the number of environments consistently improves relighting robustness, showing no performance plateau up to five training environments.
4.4 Limitation and discussion
While we achieve compelling relighting results, indirect illumination and subsurface scattering effects are not considered. As shown in Tab. 3, objects exhibiting strong shadows (e.g., hotdog, coffee) or translucent materials (e.g., ship) represent common challenging cases. Over-exposed regions also pose inherent challenges for opacity estimation to all Gaussian based relighting methods, where clipping highlights are incorrectly interpreted as transparent against white splatting backgrounds.
Table 5: Ablation on the number of environments used in training. Since up to $5$ environments are used, relighting results are reported on the remaining 3 unseen scenes: sunset, fireplace, night.
# Envs # Views Cam Poses PSNR SSIM LPIPS
1 100 identical 23.98 0.906 0.084
1 200 extra 24.07 0.908 0.081
2 200 identical 26.14 0.927 0.066
2 200 extra 25.52 0.930 0.069
3 300 identical 26.35 0.930 0.064
4 400 identical 26.68 0.931 0.062
5 500 identical 27.36 0.936 0.060
5 Conclusion
In this paper, we proposed ReCap, leveraging internal photometric consistency to address the albedo-lighting ambiguity limiting existing Gaussian relighting methods. With cross-environment captures, we explicitly model light-dependent appearance with independent learnable lighting representations that share a common set of material attributes. Combined with an optimization-friendly shading function and physically appropriate post-processing during training, ReCap demonstrates realistic relighting quality with truthful tones across diverse scenes.

### v2

@@
 Abstract
+11footnotetext: Corresponding author
 Accurate 3D objects relighting in diverse unseen environments is crucial for realistic virtual object placement.
@@
 Such coherence enables physically sound lighting reconstruction and robust material estimation — both essential for accurate relighting.
-Together with a streamlined shading function and effective post-processing, ReCap outperforms the leading competitor by $3.4$ dB in PSNR on an expanded relighting benchmark.
+Together with a streamlined shading function and effective post-processing, ReCap outperforms all leading competitors on an expanded relighting benchmark.
 Figure 1: ReCap shows a significant advantage in reconstructing environment maps with accurate tones and color fidelity.
 Both qualitative and quantitative assessments show that ReCap achieves more realistic and consistent relighting results under a range of unseen lighting conditions.
-See the project page.
 1 Introduction
@@
 Inspired by photometric appearance modeling [33, 25, 45] which estimates surface properties from object appearances under varied lighting, we introduce additional photometric supervision to address the albedo-lighting ambiguity.
-While traditional approaches rely on controlled lighting setups like light stages [11] and collocated lights [42] to provide supervision through known light directions and/or intensities, such controlled capture is impractical for real-world applications.
+While traditional approaches rely on controlled lighting setups like light stages [11] and collocated lights [44] to provide supervision through known light directions and/or intensities, they require dedicated hardwares.
 Instead, we propose ReCap to leverage object captures across unknown lighting conditions, modeling light-dependent appearances with multiple environment maps that share a common Gaussian model.
 Conceptually, this resembles multi-task learning, where the learned environment maps act as task heads querying a shared material representation for varied object appearances.
-In this case, the “querying” is done by the physically-based shading function, which facilitates accurate separation of material and lighting.
-This joint optimization fosters robust material estimation across diverse environments while maintaining internal consistency and coherence, simultaneously supporting precise lighting reconstruction.
+In this case, the “querying" is done by the physically-based shading function, which facilitates the separation of material and lighting.
+This joint optimization promotes internal consistency when accounting for varied object appearances across diverse environments, simultaneously supporting more accurate lighting reconstruction.
 To streamline the joint optimization, we introduce a generalized shading function based on the split-sum approximation [21], which eliminates a material parameter that introduces ambiguity in inverse rendering.
@@
 1, experiments on the expanded benchmark confirm the effectiveness of ReCap in producing robust and more realistic relighting results.
-Codes and datasets will be made publicly available here.
+Codes and datasets are available here.
 To summarize, the contribution of this work includes
@@
 Given $k$ sets of object appearances from unknown lighting conditions as input, $k$ learnable environment maps are instantiated.
-Gaussian color is pre-computed according to the shading function in the world space based on environment queries and material properties.
+Gaussian color is computed according to the shading function in the world space based on environment queries and material properties.
 2D images are rasterized with standard Gaussian splatting and used for loss computation.
@@
 Several NeRF-based relighting methods have taken advantage of such captures.
-NeRV [37] requires multiple known lighting environments, while TensoIR [19] internalizes lighting as an additional dimension in their factorized tensorial representation.
+NeRV [39] requires multiple known lighting environments.
 NeRD [7] models lighting as spherical Gaussians and optimizes separate lighting for every input image.
-In contrast, we use object captures from unknown lighting conditions, with learnable lighting representations optimized independently for each scene as environment maps.
+TensoIR [19] encodes lighting as an additional dimension in its factorized tensor representation and relies on ground-truth albedo scaling for accurate relighting.
+In contrast, we use object captures under unknown lighting, optimizing a learnable light map separately for each scene and independently from the Gaussian model, without requiring ground-truth albedo for relighting.
 3 Method
@@
 Substitute Eq. 5 into Eq. 3 and apply the split-sum approximation as described in  [21], the shading function can be written as
-$\begin{split}L_{out}(\mathbf{x},\mathbf{v})&=\underbrace{E_{d}(\mathbf{n})(1-m)\mathbf{b}}_{diffuse}\\ &+\underbrace{E_{s}(\mathbf{n},\mathbf{v})\left[F_{0}\beta_{1}(r,\mathbf{n},\mathbf{v})+\beta_{2}(r,\mathbf{n},\mathbf{v})\right]}_{specular}\\ \end{split}$ (6)
+$\begin{split}L_{out}(\mathbf{x},\mathbf{v})&=\underbrace{E_{d}(\mathbf{n})(1-m)\mathbf{b}}_{diffuse}\\ &+\underbrace{E_{s}(\mathbf{n},\mathbf{v})\left[F_{0}\beta_{1}(r,\mathbf{n},\mathbf{v})+\beta_{2}(r,\mathbf{n},\mathbf{v})\right]}_{specular},\\ \end{split}$ (6)
 where $E_{d}$ and $E_{s}$ are the pre-filtered environment maps for diffuse and specular reflectance, $\beta_{1}$ and $\beta_{2}$ are pre-calculated BRDF look-ups, $F_{0}=m\mathbf{b}+(1-m)s$ is the effective reflectance.
@@
 4.3, it also requires additional effort to capture real objects across multiple scenes.
-As a proof of concept, we limit our investigation to a dual-environment setup for the remainder of the text.
+As a proof of concept, we limit our investigation to a dual-environment setup unless mentioned otherwise.
 Figure 4: The same object position exhibit view-dependent and light-dependent pixel color, the later is accounted for by querying corresponding learnable environments.
-3.4 Physically Appropriate Post-processing
-Standard HDR maps typically represent radiance in linear RGB space when used as lighting sources in rendering [10].
-To properly apply them for relighting, the learned lighting values need a clear physical interpretation.
-Current methods make diverse design choices during training, which jointly determine the interpretation of the learned environment maps and how new HDR maps should be processed.
-For example, if the range of the learned values is limited between $[0,1]$ and gamma correction is applied after the shading equation, the learned environment map is interpreted as being in a linear low dynamic range (LDR) space.
+3.4 Post-shading Processing
+Standard HDR maps represent radiance in linear RGB space when used as lighting sources in rendering [10].
+To apply them effectively for relighting, the learned lighting values need a clear physical interpretation.
+Current methods adopt diverse post-shading strategies during training, influencing both the interpretation of learned environment maps and the processing of new HDR maps.
+For example, if the learned values are constrained to $[0,1]$ and gamma correction is applied post-shading, the environment map is effectively treated as being in a linear low dynamic range (LDR) space.
 In Tab.
-1, we analyze three major factors that impact relighting performance.
-Gamma correction is found to be essential for relighting, which is expected, since the shading function models linear light transport.
-However, it is often overlooked [15, 44], as its omission still provides good NVS results.
-Complex tone-mappers are common in graphics and are widely adopted [37], but they perform worse than simple clipping, likely due to the introduced non-linearity that hinders optimization.
-Striking a balance between relighting and NVS, we constrain our environment maps to be positive, applying simple clipping followed by gamma correction.
-This ensures a linear HDR lighting space, where new environment maps can be applied without modification.
-Together with cross-environment decoupling, our learned environment maps are closer to the actual HDR maps in distribution.
-This removes the need for normalizing relit images [26] to match the average colors of ground-truth images (which are not available in practice) when comparing relighting results.
+1, we analyze three major factors that affects relighting.
+Gamma correction is essential for relighting, as expected, since the shading function models linear light transport.
+However, it is often overlooked [15, 46], because omitting it still yields good NVS results.
+While previous work also consider complex tone-mappers [39] from graphics, they perform worse than simple clipping, likely due to the introduced non-linearity hindering optimization.
+To balance relighting and NVS, we constrain environment maps to be positive and apply simple clipping followed by gamma correction.
+This implies a linear HDR lighting space where new environment maps can be used without modification, removing the need for normalizing relit images [27] or learned albedo [19] to match the ground-truth (unavailable in practice) when when evaluating relighting.
+Combined with cross-environment decoupling, this essentially improves relighting by (i) allowing Gaussians to push light-dependent appearances into learnable “sinks" as pseudo light maps, and (ii) regulating these “sinks" with HDR processing and PBR, enabling them to approximate real ambient light in value distribution.
 Table 1: Relighting and NVS performance (PSNR) of various design choices.
@@
 Cross-environment supervision provides more robust normal estimation and correct highlight shapes.
-Table 2: Relighting results across different environment maps.
-ReCap trains on 100 views each from “courtyard” and “bridge”, while other methods use 200 views from “bridge” only.
-The * symbol indicates learned environment, equivalent to NVS.
-Note that 3DGS-DR’s good performance on “bridge” reflects its tendency to regurgitate training views.
-Best results are highlighted in bold.
-Unseen Envs Training Envs NVS
-night snow sunset interior fireplace forest average courtyard bridge bridge*
+Table 2: Relighting and NVS results.
+Methods optimized over single-env ( ) and dual-env ( ) inputs differ in the choice of training split B.
+“GT scaling" refers to the use of ground truth albedo to re-scale learned albedo for relighting.
+Best and second best results are highlighted in bold and italic respectively.
+use GT scaling Training Envs Unseen Relights Seen Relights NVS
+split A split B night snow sunset interior fireplace forest Avg. bridge courtyard bridge* courtyard*
 […]
-GShader [18] 22.33 18.14 19.26 21.87 21.03 23.04 20.95 20.68 26.04 32.92
-3DGS-DR [44] 20.88 18.92 23.53 21.00 19.52 25.19 21.51 20.58 33.19 35.50
-GS-IR [25] 24.87 21.95 21.15 20.18 24.27 20.47 22.15 20.02 20.72 29.95
-R3DG [15] 23.20 22.05 22.93 19.65 22.62 22.26 22.12 20.22 22.25 30.44
-ReCap (ours) 26.68 24.56 26.56 24.34 25.51 25.45 25.52 25.51 26.97 32.03
+3DGS-DR[46] - bridge bridge 20.18 18.99 23.19 21.33 18.76 24.90 21.22 33.18 - 35.59 -
+GS-IR[26] - bridge bridge 25.56 22.44 21.22 19.46 25.64 21.10 22.57 22.01 - 31.22 -
+R3DG[15] - bridge bridge 23.71 22.09 22.84 20.09 23.08 22.19 22.33 22.32 - 30.61 -
+GShader[18] - bridge bridge 21.48 17.61 18.93 21.89 20.18 22.70 20.47 26.03 - 33.05 -
+ReCap (ours) - bridge bridge 25.31 21.76 23.20 21.66 24.52 23.15 23.27 25.30 - 33.40 -
+3DGS-DR[46] - bridge courtyard 20.53 20.25 24.56 21.71 19.41 24.25 21.78 24.89 23.92 24.84 23.57
+GS-IR[26] - bridge courtyard 26.22 22.03 20.85 19.18 25.90 20.52 22.45 20.78 20.96 24.92 23.16
+R3DG[15] - bridge courtyard 23.91 21.78 22.55 19.86 23.38 21.74 22.21 20.98 20.84 25.26 22.68
+GShader[18] - bridge courtyard 22.73 18.20 19.96 22.11 20.97 23.02 21.17 23.34 21.75 24.96 22.84
+TensoIR[19] bridge courtyard 27.22 22.15 24.31 23.12 25.35 24.80 24.49 24.50 23.82 29.13 27.46
+TensoIRno scale[19] - bridge courtyard 26.24 20.11 22.75 21.79 24.45 23.29 23.11 23.50 22.47 29.13 27.46
+GShader[18] + ours - bridge courtyard 23.52 17.69 19.95 23.22 21.75 24.50 21.77 24.45 24.32 31.13 29.30
+ReCap (ours) - bridge courtyard 27.62 24.64 26.33 24.40 26.52 25.38 25.82 26.95 25.52 32.23 30.76
 […]
-GShader [18] 0.899 0.889 0.900 0.907 0.869 0.907 0.895 0.893 0.950 0.969
-3DGS-DR [44] 0.892 0.897 0.931 0.909 0.858 0.910 0.900 0.894 0.971 0.978
-GS-IR [25] 0.900 0.896 0.892 0.878 0.876 0.863 0.884 0.871 0.879 0.938
-R3DG [15] 0.889 0.914 0.920 0.865 0.867 0.879 0.889 0.871 0.879 0.959
-ReCap (ours) 0.932 0.938 0.951 0.929 0.903 0.927 0.930 0.942 0.951 0.965
+3DGS-DR[46] - bridge bridge 0.882 0.894 0.925 0.907 0.843 0.906 0.893 0.971 - 0.978 -
+GS-IR[26] - bridge bridge 0.902 0.904 0.890 0.861 0.884 0.867 0.885 0.898 - 0.952 -
+R3DG[15] - bridge bridge 0.889 0.913 0.918 0.866 0.866 0.877 0.888 0.879 - 0.959 -
+GShader[18] - bridge bridge 0.889 0.887 0.896 0.905 0.854 0.904 0.889 0.949 - 0.968 -
+ReCap (ours) - bridge bridge 0.911 0.918 0.926 0.907 0.883 0.906 0.909 0.946 - 0.970 -
+3DGS-DR[46] - bridge courtyard 0.885 0.902 0.934 0.907 0.846 0.903 0.896 0.930 0.927 0.926 0.924
+GS-IR[26] - bridge courtyard 0.906 0.893 0.887 0.858 0.885 0.862 0.882 0.881 0.879 0.915 0.905
+R3DG[15] - bridge courtyard 0.896 0.912 0.922 0.866 0.870 0.873 0.890 0.859 0.888 0.932 0.919
+GShader[18] - bridge courtyard 0.904 0.891 0.908 0.906 0.862 0.905 0.896 0.925 0.913 0.928 0.922
+TensoIR[19] bridge courtyard 0.908 0.861 0.891 0.870 0.886 0.888 0.884 0.893 0.883 0.962 0.957
+TensoIRno scale[19] - bridge courtyard 0.910 0.861 0.893 0.871 0.889 0.889 0.885 0.895 0.884 0.962 0.957
+GShader[18] + ours - bridge courtyard 0.915 0.893 0.915 0.919 0.874 0.922 0.906 0.937 0.938 0.959 0.956
+ReCap (ours) - bridge courtyard 0.935 0.938 0.951 0.929 0.903 0.926 0.930 0.951 0.943 0.966 0.963
 […]
-GShader [18] 0.086 0.115 0.095 0.091 0.106 0.089 0.102 0.097 0.061 0.043
-3DGS-DR [44] 0.077 0.089 0.070 0.086 0.101 0.081 0.084 0.095 0.040 0.033
-GS-IR [25] 0.104 0.102 0.105 0.112 0.112 0.114 0.108 0.122 0.113 0.071
-R3DG [15] 0.104 0.088 0.081 0.126 0.113 0.103 0.108 0.113 0.113 0.053
-ReCap (ours) 0.062 0.070 0.057 0.076 0.076 0.071 0.069 0.065 0.058 0.047
+3DGS-DR[46] - bridge bridge 0.081 0.091 0.073 0.085 0.108 0.083 0.087 0.041 - 0.034 -
+GS-IR[26] - bridge bridge 0.099 0.094 0.101 0.119 0.108 0.109 0.105 0.099 - 0.065 -
+R3DG[15] - bridge bridge 0.103 0.086 0.082 0.124 0.114 0.104 0.102 0.113 - 0.053 -
+GShader[18] - bridge bridge 0.091 0.116 0.098 0.094 0.113 0.091 0.100 0.062 - 0.044 -
+ReCap (ours) - bridge bridge 0.077 0.080 0.073 0.089 0.092 0.084 0.083 0.061 - 0.042 -
+3DGS-DR[46] - bridge courtyard 0.084 0.096 0.078 0.096 0.115 0.094 0.094 0.078 0.080 0.082 0.085
+GS-IR[26] - bridge courtyard 0.101 0.107 0.107 0.127 0.112 0.116 0.112 0.115 0.119 0.099 0.104
+R3DG[15] - bridge courtyard 0.101 0.092 0.083 0.129 0.117 0.110 0.105 0.127 0.106 0.078 0.089
+GShader[18] - bridge courtyard 0.087 0.115 0.095 0.101 0.114 0.097 0.101 0.086 0.094 0.080 0.086
+TensoIR[19] bridge courtyard 0.131 0.155 0.134 0.138 0.138 0.127 0.137 0.128 0.133 0.093 0.098
+TensoIRno scale[19] - bridge courtyard 0.133 0.159 0.136 0.141 0.138 0.131 0.140 0.130 0.137 0.093 0.098
+GShader[18] + ours - bridge courtyard 0.074 0.113 0.084 0.089 0.102 0.080 0.090 0.073 0.073 0.055 0.059
+ReCap (ours) - bridge courtyard 0.059 0.069 0.057 0.077 0.075 0.071 0.068 0.058 0.064 0.047 0.051
 4 Experiments
@@
 Each scene includes 200 training views and 200 test views, with identical camera poses across scenes.
-For a consistent relighting comparison, all methods are trained with captures from a single environment, except ReCap, which additionally uses a second environment.
-To ensure fairness in the number of training views, all methods use 200 views from the primary training environment, while ReCap uses 100 views from each of the two training environments.
-Relighting performance will be judged on unseen environments.
+To eliminate biases from training view poses, the training views are divided into two splits, A and B, each containing 100 views.
+In the single-environment (single-env) setup, both splits come from the same environment, whereas in the dual-environment (dual-env) setup, split B is taken from a different environment.
+Relighting performance is judged on unseen environments.
 Baseline and Metrics.
-We compare our method against the following baselines: a) GShader [18]: A Gasussian rendering method utilizing a customized shading function with a residual color term.
-b) GS-IR [25]: An inverse rendering approach using view-space shading with a microfacet BRDF model and a baked occlusion map.
-c) 3DGS-DR [44]: A Gasussian rendering method leveraging a customized shading function and normal propagation in view-space.
-d) R3DG [15]: A Gaussian relighting technique that incorporates ray tracing and a microfacet BRDF model.
+We compare with the following baselines: (a) GShader [18]: a Gaussian rendering method utilizing a customized shading function with a residual color term; (b) GS-IR [26]: an inverse rendering approach using view-space shading with a microfacet BRDF model and a baked occlusion map; (c) 3DGS-DR [46]: a Gasussian rendering method targeting reflective surfaces with a customized shading function and normal propagation; (d) R3DG [15]: a Gaussian relighting technique that incorporates ray tracing and a microfacet BRDF model; and (e) TensoIR [19]: top-performing NeRF-based method in novel view synthesis and relighting.
 Following these works, rendering and relighting quality are evaluated using PSNR, SSIM [42], and LPIPS [49].
@@
 All experiments are conducted on an Nvidia RTX 3090 graphics card.
-We optimize our models using the Adam optimizer over 30,000 iterations.
-All methods are retrained on our dataset using their official repositories with recommended settings, results are reported from the retrained models.
-Since GShader [18] and 3DGS-DR [44] do not provide official relighting code, we pre-processed the HDR maps according to their design choices, with details in the supplementary material.
-Figure 6: Qualitative comparison of relighting results across different environment maps.
-Specular objects provide more information for lighting reconstruction, leading to better albedo-lighting decoupling hence more truthful relighting results.
+Our models are optimized using Adam for 30,000 iterations.
+All other methods are retrained on our dataset using their official repositories and recommended settings, with results reported from the retrained models.
+As GShader [18] and 3DGS-DR [46] lack official relighting code, we pre-processed HDR maps following their design choices (details in the supplementary).
+For TensoIR [19], which uses per-object hyperparameters, we tested all provided settings and report the best results.
+Figure 6: Qualitative comparison of NVS and relighting results.
+Upper: toaster, lower: chair.
+TensoIR produces accurate relighting with GT albedo scaling for chair, but suffers from tone-shift without it.
+For our method, specular surfaces in toaster provide more information for lighting reconstruction, leading to better albedo-lighting decoupling hence more truthful relighting results than mostly diffuse chair.
 Refer to the supplementary for more visuals.
-Table 3: Object-wise relighting performance.
-The average is reported over unseen environment maps.
-ReCap shows the best overall performance for both diffuse and specular objects, while performing comparably on objects with shadows or translucent materials.
-Best results are highlighted in bold.
-Relit NeRF Synthetic Relit Shiny Blender
-chair drums ficus hotdog lego ship material mic helmet coffee musclecar toaster teapot
-[…]
-GShader [18] $20.83$ $19.69$ $25.56$ $18.80$ $18.85$ $17.10$ $20.03$ $27.86$ $19.07$ $\mathbf{20.28}$ $21.51$ $18.55$ $24.14$
-3DGS-DR [44] $21.25$ $21.82$ $25.78$ $19.34$ $24.37$ $18.98$ $20.01$ $28.02$ $20.33$ $18.56$ $19.47$ $17.35$ $24.34$
-GS-IR [25] $21.87$ $22.37$ $26.75$ $21.74$ $19.89$ $\mathbf{21.18}$ $22.26$ $27.25$ $19.31$ $18.60$ $23.39$ 16.79 26.52
-R3DG [15] 23.67 21.23 25.60 24.38 22.21 20.50 21.55 27.67 21.17 16.27 22.74 16.66 23.89
-ReCap (ours) 26.63 24.77 28.24 23.25 24.44 20.12 25.83 31.44 23.55 19.78 29.14 25.05 29.49
-[…]
-GShader [18] 0.914 0.891 0.951 0.885 0.866 0.774 0.871 0.968 0.889 0.931 0.898 0.824 0.973
-3DGS-DR [44] 0.920 0.911 0.958 0.882 0.926 0.771 0.875 0.970 0.885 0.921 0.881 0.824 0.971
-GS-IR [25] 0.905 0.901 0.951 0.886 0.856 0.783 0.873 0.958 0.850 0.896 0.896 0.774 0.967
-R3DG [15] 0.924 0.895 0.952 0.904 0.888 0.775 0.864 0.962 0.865 0.877 0.900 0.784 0.968
-ReCap (ours) 0.951 0.937 0.972 0.912 0.918 0.817 0.921 0.978 0.920 0.928 0.944 0.910 0.983
-[…]
-GShader [18] 0.078 0.068 0.038 0.133 0.099 0.187 0.089 0.024 0.134 0.142 0.073 0.157 0.038
-3DGS-DR [44] 0.055 0.056 0.029 0.110 0.055 0.173 0.081 0.020 0.116 0.144 0.075 0.143 0.033
-GS-IR [25] 0.068 0.069 0.040 0.133 0.096 0.189 0.093 0.032 0.178 0.164 0.083 0.220 0.041
-R3DG [15] 0.060 0.068 0.037 0.105 0.084 0.188 0.103 0.033 0.171 0.157 0.078 0.210 0.038
-ReCap (ours) 0.042 0.042 0.023 0.091 0.065 0.167 0.063 0.016 0.094 0.128 0.047 0.088 0.027
 4.2 Comparison with Previous Methods
-Scene-wise Comparison.
-ReCap consistently outperforms other methods on all three metrics across various unseen environments, as shown in  Tab.
-2, while maintaining comparable NVS performance.
-As further supported by the visual comparisons in Fig.
-6 where ReCap demonstrates noticeable perceptual improvements over other methods, this indicates advantageous relighting quality and fidelity without sacrificing NVS performance.
-A notable observation across all methods is the large performance gap between NVS using the learned environment map and relighting using the ground truth HDR map (bridge* vs bridge).
-This highlights the albedo-lighting ambiguity, where the learned lighting representation diverges from ground truth but still works effectively with the learned materials for NVS.
-Among all methods, 3DGS-DR [44] shows the smallest gap, nonetheless its performance on unseen environments remains limited.
-As shown by the second row in Fig.
-6, this is because that 3DGS-DR tends to preserve a consistent tone across different environments, closely matching the training views and adapting only the specular reflections based on new lighting.
-Object-wise Comparison.
-While scene-wise comparison confirms our method’s ability to generalize across different lighting conditions, a more granular evaluation of how effectively it adapts to diverse object attributes is presented in Tab.
-3.
-Across various test objects featuring diffuse and specular surfaces, ReCap shows the best overall performance.
-Especially with more specular objects (e.g., toaster, musclecar) where other methods find challenging, ReCap shows a unique advantage in decoupling rich reflected lighting details for accurate relighting.
+Single-env Setup.
+We compare single-env performance in the first group of Tab.
+2.
+The substantial difference between NVS using the learned environment map and relighting using the ground truth HDR map (bridge* vs.
+bridge) highlights the albedo-lighting ambiguity: the learned lighting representation diverges from ground truth but still works effectively with the learned materials for NVS.
+While 3DGS-DR [46] has the smallest gap, its ability to generalize to unseen relighting conditions remains limited.
+As shown in the second column of Fig.
+6, this is because it tends to reproduce training views, preserving the overall tone and adapting only the specular reflections to new lighting.
+Dual-env Setup.
+The second group of Tab.
+2 presents results from directly optimizing existing Gaussian models over dual-env inputs while maintaining a single learned lighting representation.
+This leads to mixed relighting performance and significant NVS artifacts since the learned environment is no longer well-defined.
+Therefore, we continue our analysis with their single-env results for the remainder of the discussion.
+In the third group, we compare against TensoIR [19] with their multi-light setup and further augment GShader [18] with the proposed multi-light querying.
+Since TensoIR [19] uses GT albedo scaling in their official code, we additionally report their scale-free performance for clearer comparison.
+Qualitative Comparison.
+Fig.
+6 compares NVS and relighting performance across single-env Gaussian relighting methods, two dual-env TensoIR [19] variants, and dual-env ReCap.
+Our ReCap achieves noticeable perceptual improvements over methods that do not rely on GT for relighting.
+TensoIR [19] produces accurate relighting with GT albedo scaling, but suffers from obvious tone-shift without it (e.g., chair) and struggles with metallic objects (e.g., toaster) due to their dielectric assumption.
 Reconstructed Environments.
@@
 7, we display the reconstructed environment maps for objects with varying level of specularity.
-When the object is mostly diffuse, it’s challenging to disentangle lighting effects from the object’s intrinsic properties due to insufficient lighting cues.
+For predominantly diffuse objects, disentangling lighting from intrinsic properties is challenging due to limited lighting cues.
+Notice how our learned light map for chair retains object color, highlighting this challenge.
 Figure 7: Learned environment maps for different objects. As objects become more diffuse, reconstructions are less accurate.
@@
 8, we compare the average relighting performance across unseen scenes in  Tab.
-4.
+3.
 The new shading function yields the most substantial improvement by offering a more flexible material representation, with the two additional regularization terms also contributing positively to the final results.
@@
 Extra training environments.
-Up to this point, we have focused on a dual-environment setup.
-In this experiment, we examine how extra photometric supervision from additional training environments affects the relighting results.
+So far, we have focused on dual-env setups.
+Here, we examine how extra photometric supervision from additional training environments affects the relighting.
 Due to the constraint of 200 training views per scene, we allocate 100 identical views per environment as we include more environments for training.
-To provide a basis for comparison with existing results, we additionally compare the use of identical views versus extra views.
+To align with existing results, we also compare the use of identical views versus extra views.
 Results in Tab.
-5 highlight three key findings: 1) more views from additional environment yields better results than adding more views within the same environment; 2) using identical views across environments provides stronger feature decoupling; and 3) increasing the number of environments consistently improves relighting robustness, showing no performance plateau up to five training environments.
+4 reveal three key findings: (i) Adding identical views from an extra environment is better than adding extra unique views within the same environment (row 2 vs.
+3); (ii) Although identical views across environments theoretically provides stronger decoupling, the benefit of extra unique views is more pronounced in practice (row 3 vs.
+4).
+It is likely that extra views already provide sufficient decoupling, especially with perfect pose calibration in synthetic datasets; and (iii) Increasing the number of environments consistently enhances relighting performance, with no observed plateau up to five.
+Table 4: Ablation on the number of environments used in training. Since up to $5$ environments are used, relighting results are reported on the remaining 3 unseen scenes: sunset, fireplace, night.
+# Envs # Views # Unique Views Cam Poses PSNR SSIM LPIPS
+1 100 100 identical 23.98 0.906 0.084
+1 200 200 extra 24.07 0.908 0.081
+2 200 100 identical 26.14 0.927 0.066
+2 200 200 extra 26.25 0.928 0.065
+3 300 100 identical 26.35 0.930 0.064
+4 400 100 identical 26.68 0.931 0.062
+5 500 100 identical 27.36 0.936 0.060
+Application on Real Captures.
+ReCap requires multiple sets of camera poses to align to a common coordinate.
+In Fig.
+8, we demonstrate its practicality with two real-life examples from StanfordORB [24], which captures the same object in different environments and provides COLMAP-estimated poses for each environment.
+They also provide sparse cross-environment alignment pairs computed by SuperGlue [38].
+We use these pairs to estimate the view transform matrix and account for scale ambiguity.
+While exact calibration is non-trivial, sufficiently accurate pose calibration proves to be feasible for the proposed application.
+Figure 8: Relighting and NVS examples on StanfordORB: gnome, teapot.
 4.4 Limitation and discussion
 While we achieve compelling relighting results, indirect illumination and subsurface scattering effects are not considered.
-As shown in Tab.
-3, objects exhibiting strong shadows (e.g., hotdog, coffee) or translucent materials (e.g., ship) represent common challenging cases.
-Over-exposed regions also pose inherent challenges for opacity estimation to all Gaussian based relighting methods, where clipping highlights are incorrectly interpreted as transparent against white splatting backgrounds.
-Table 5: Ablation on the number of environments used in training. Since up to $5$ environments are used, relighting results are reported on the remaining 3 unseen scenes: sunset, fireplace, night.
-# Envs # Views Cam Poses PSNR SSIM LPIPS
-1 100 identical 23.98 0.906 0.084
-1 200 extra 24.07 0.908 0.081
-2 200 identical 26.14 0.927 0.066
-2 200 extra 25.52 0.930 0.069
-3 300 identical 26.35 0.930 0.064
-4 400 identical 26.68 0.931 0.062
-5 500 identical 27.36 0.936 0.060
+Over-exposed regions also present inherent challenges for opacity estimation in all Gaussian-based relighting methods, where clipped highlights are incorrectly interpreted as transparent against white splatting backgrounds.
+For easier pose calibration in practice, semi-controlled capturing, such as rotating the object within the same environment at a known angle, may be considered.
 5 Conclusion
@@
 Combined with an optimization-friendly shading function and physically appropriate post-processing during training, ReCap demonstrates realistic relighting quality with truthful tones across diverse scenes.
+Acknowledgment. This research was supported by the Alexander von Humboldt Foundation.

### v3

@@
 Acknowledgment. This research was supported by the Alexander von Humboldt Foundation.
+Supplementary Material
+In the supplementary material, we first provide detailed explanations of the HDR processing steps in  Sec.
+A.
+Next, we illustrate the effects of the proposed shading function in Sec.
+B.
+Additionally, object-wise comparison and extended qualitative comparisons are presented in Sec.
+C and Sec.
+D.
+A HDR Processing for Relighting
+In Sec.
+3.4, we discussed how various design choices influence the interpretation of the learned environment maps and how novel HDR maps should be processed to align with this interpretation for relighting.
+Fig.
+A summarizes the four options, with the red dotted block indicating the processing applied to standard HDR maps to obtain the relighting results presented in Tab.
+1 of the main text.
+While omitting gamma correction during training is theoretically inconsistent with a linear light transport, it is included for completeness and to reflect common practice where it is often overlooked.
+Figure A: Design choices and relevant implication on the learned environment maps. During relighting (highlighted by the red dotted block), the input HDR map are processed to match the interpretation.
+Since GShader [18] and 3DGS-DR [46] do not provide official relighting code, Sec. A.1 and Sec. A.2 explain the processing applied to obtain their relighting results.
+A.1 GShader
+GShader [18] constrains its learned environment maps to the range $[0,1]$ and adopts a $y$-up coordinate convention for light querying.
+As illustrated in Fig.
+B, we perform coordinate transformations and clipping on latitude-longitude HDR environment maps for relighting.
+Without clipping, the rendered object appears overexposed.
+Figure B: Environment map processing for GShader.
+A.2 3DGS-DR
+The learned environments in 3DGS-DR [46] are not range-limited and adopts a $y$-down coordinate system.
+Since the official implementation visualizes learned environment maps with a sigmoid function, we use an inverse sigmoid to adjust the lighting values.
+As shown in Fig.
+C, we apply coordinate transformation, clipping and inverse sigmoid on the HDR maps for relighting.
+Without the adjustment, the relighting output appears dull, and the highlight pattern also shifts incorrectly (from the upper surface to the front surface) due to the mismatched coordinate system.
+Figure C: Environment map processing for 3DGS-DR.
+B Visual Ablation of the Shading Function
+The proposed shading function offers a more flexible material representation, which proves particularly advantageous for specular surfaces as shown in Fig.
+D.
+Although relighting differences are less pronounced for diffuse objects, training by the original shading function tends to misclassify non-metallic surfaces, such as the ficus leaves.
+Figure D: Comparison of the original and the proposed shading function.
+Table E: Object-wise relighting performance.
+Methods optimized over single-env ( ) and dual-env ( ) inputs differ in the choice of training split B.
+Best and second best results are highlighted in bold and italic respectively.
+Relit NeRF Synthetic Relit Shiny Blender
+chair drums ficus hotdog lego ship material mic helmet coffee musclecar toaster teapot
+[…]
+3DGS-DR[46] 21.25 21.82 25.78 19.34 21.38 19.06 20.01 26.42 20.33 18.09 19.62 17.35 25.46
+GS-IR[26] 20.56 22.53 27.45 22.22 24.00 21.79 21.38 27.58 19.63 20.40 22.93 16.51 26.44
+R3DG[15] 23.67 21.23 25.60 24.38 24.13 20.53 21.55 26.77 21.17 17.66 22.47 16.66 24.54
+GShader[18] 20.83 19.69 25.56 18.80 16.42 17.37 20.03 26.16 19.07 18.35 21.19 18.55 24.03
+ReCap (ours) 21.49 24.02 24.82 22.43 23.22 20.08 22.69 29.70 23.41 18.38 25.74 18.89 27.59
+3DGS-DR[46] 21.86 22.09 25.55 21.13 22.71 20.41 20.73 26.60 18.54 20.44 21.66 15.14 26.33
+GS-IR[26] 21.18 22.70 27.21 21.74 24.07 21.51 21.80 26.84 19.41 20.27 22.74 16.05 26.34
+R3DG[15] 24.08 21.25 25.17 23.57 24.50 18.34 22.17 25.76 20.74 18.66 22.09 17.01 25.33
+GShader[18] 21.33 21.11 25.71 18.70 18.47 18.98 20.98 27.17 19.51 18.28 22.63 18.06 24.24
+TensoIR[19] 26.32 20.77 24.29 26.06 27.08 21.01 20.69 29.77 20.11 25.13 25.83 15.43 35.89
+TensoIRno scale[19] 20.25 22.99 24.28 24.50 28.23 21.98 19.96 28.40 18.76 23.66 24.59 15.02 27.76
+GShader[18] + ours 23.25 21.06 26.63 20.24 19.17 18.80 19.97 27.18 19.81 20.34 21.84 20.53 24.20
+ReCap (ours) 26.63 24.77 28.24 23.25 24.04 20.12 25.83 31.31 23.55 22.17 28.99 25.05 31.65
+[…]
+3DGS-DR[46] 0.920 0.911 0.958 0.882 0.867 0.771 0.875 0.956 0.885 0.903 0.882 0.824 0.974
+GS-IR[26] 0.897 0.901 0.956 0.889 0.874 0.790 0.864 0.954 0.849 0.896 0.895 0.767 0.969
+R3DG[15] 0.924 0.895 0.952 0.904 0.880 0.774 0.864 0.955 0.865 0.881 0.899 0.784 0.969
+GShader[18] 0.914 0.891 0.951 0.885 0.821 0.777 0.871 0.956 0.889 0.908 0.896 0.824 0.973
+ReCap (ours) 0.922 0.928 0.961 0.900 0.887 0.801 0.889 0.970 0.909 0.910 0.928 0.824 0.982
+3DGS-DR[46] 0.925 0.916 0.957 0.897 0.884 0.804 0.876 0.957 0.862 0.919 0.896 0.781 0.976
+GS-IR[26] 0.900 0.901 0.954 0.885 0.868 0.793 0.861 0.950 0.848 0.894 0.888 0.756 0.966
+R3DG[15] 0.931 0.897 0.951 0.904 0.883 0.761 0.871 0.952 0.867 0.891 0.899 0.792 0.972
+GShader[18] 0.920 0.901 0.954 0.892 0.843 0.794 0.878 0.961 0.889 0.909 0.911 0.820 0.974
+TensoIR[19] 0.939 0.877 0.946 0.911 0.922 0.777 0.862 0.970 0.825 0.891 0.910 0.684 0.978
+TensoIRno scale[19] 0.920 0.893 0.946 0.912 0.925 0.800 0.870 0.967 0.818 0.892 0.901 0.688 0.977
+GShader[18] + ours 0.935 0.905 0.960 0.900 0.855 0.795 0.888 0.963 0.903 0.927 0.906 0.869 0.975
+ReCap (ours) 0.951 0.937 0.972 0.912 0.908 0.815 0.921 0.975 0.920 0.940 0.943 0.910 0.988
+[…]
+3DGS-DR[46] 0.055 0.056 0.029 0.110 0.088 0.173 0.081 0.025 0.116 0.149 0.075 0.143 0.030
+GS-IR[26] 0.072 0.069 0.036 0.121 0.090 0.185 0.094 0.032 0.176 0.152 0.080 0.221 0.037
+R3DG[15] 0.060 0.068 0.037 0.105 0.088 0.188 0.103 0.035 0.171 0.152 0.078 0.210 0.035
+GShader[18] 0.078 0.068 0.038 0.133 0.119 0.192 0.089 0.028 0.134 0.158 0.075 0.157 0.037
+ReCap (ours) 0.060 0.048 0.028 0.103 0.081 0.162 0.082 0.019 0.118 0.140 0.057 0.151 0.023
+3DGS-DR[46] 0.057 0.055 0.032 0.102 0.085 0.187 0.087 0.028 0.148 0.150 0.073 0.185 0.031
+GS-IR[26] 0.074 0.070 0.040 0.131 0.099 0.196 0.102 0.037 0.183 0.161 0.085 0.235 0.040
+R3DG[15] 0.060 0.067 0.039 0.110 0.091 0.206 0.104 0.038 0.170 0.160 0.078 0.212 0.037
+GShader[18] 0.076 0.064 0.038 0.130 0.106 0.219 0.090 0.027 0.135 0.164 0.068 0.164 0.038
+TensoIR[19] 0.075 0.122 0.067 0.135 0.096 0.221 0.134 0.043 0.256 0.164 0.116 0.310 0.046
+TensoIRno scale[19] 0.108 0.119 0.067 0.144 0.089 0.208 0.132 0.046 0.262 0.165 0.121 0.308 0.048
+GShader[18] + ours 0.063 0.062 0.032 0.100 0.095 0.205 0.088 0.024 0.125 0.146 0.071 0.124 0.038
+ReCap (ours) 0.042 0.042 0.023 0.091 0.068 0.165 0.063 0.016 0.094 0.120 0.048 0.088 0.021
+C Object-wise results
+We present object-wise relighting performance in Tab.
+E.
+Objects with strong self-shadowing/self-reflection (e.g., hotdog, coffee, lego as shown in Fig.
+E) are difficult for ReCap as we do not handle secondary rays.
+Ship poses a challenge for all methods with its translucent material requiring the modeling of subsurface scattering.
+D Additional Results (Fig. F to Fig. H)
+Figure E: Training views of objects with self-shadowing/self-reflection.
+Figure F: Qualitative comparison of relighting results across different environment maps. Upper: toaster; lower: helmet.
+Figure G: Qualitative comparison of relighting results across different environment maps. Upper: musclecar; lower: chair.
+Figure H: Qualitative comparison of relighting results across different environment maps. Upper: mic; lower: ficus.